# Relationship Discovery for Drug Recommendation

### v1

@@ 4.3.1 Patient Conditions Difference Capturing
Taking the example of diseases, for the analysis of differences between diseases in a clinical visit, we categorize the diseases $\mathcal{D}_{t}$ into four types $\mathcal{D}^{j}_{t}$ based on $G^{d}_{v_{t}}$: 1) Causal diseases, which initiate other diseases and represent the primary ailment; 2) Effect diseases, influenced by other diseases and indicative of secondary symptoms; 3) Middle diseases, both influencing and influenced by other diseases; 4) Independent diseases, existing without direct causal links to other diseases.
Subsequently, we introduce Dynamic Self-Adaptive Attention (DSA), designed to allocate distinct weights to different disease groups across various clinical visits. As diseases within a single medical visit are classified into multiple distinct groups, the DSA mechanism can accurately discern differences among various diseases. Furthermore, considering that the grouping of diseases dynamically changes across different visits, DSA is equally capable of effectively capturing these dynamic differences of diseases.
$\footnotesize\mathcal{D}^{j}_{t}=\text{Classify}(d_{i},G^{d}_{v_{t}}),$ (4)
$\mathbf{h}^{r}_{d_{i}}=\mathbf{h}_{d_{i}}\cdot w^{j}_{t},$ (5)
$w^{j}_{t}=\frac{\exp(\mathbf{W}\cdot\mathbf{h}^{D^{j}}_{t}+b)}{\sum_{k=1}^{4}\exp(\mathbf{W}\cdot\mathbf{h}^{D^{k}}_{t}+b)},$ (6)
@@
4.3.2 Refinement of Point-to-Point Relationship
In this stage, we refine the point-to-point relationships between diseases/procedures and medications based on causal effects. We utilize the causal effect matrices $\mathbf{M}^{dm}$ and $\mathbf{M}^{pm}$ to construct disease-medication and procedure-medication complete bipartite graphs $G^{dm}_{v_{t}}$ and $G^{pm}_{v_{t}}$ for single visit.
Taking $G^{dm}_{v_{t}}$ as an example, the nodes encompass both $\mathcal{D}_{t}$ and $\mathcal{M}_{t-1}$. By utilizing Relational Graph Convolutional Networks (RGCN) Schlichtkrull et al. (2018), we categorize edges corresponding to relationships with similar causal effects into the same type. Subsequently, we establish point-to-point relationships between diseases and medications. The representation of $d_{i}$ at layer $l$ is then updated on the point-to-point relationship using the following formula:
$\footnotesize\mathbf{h}_{d_{i}}^{l+1}=\sigma(\mathbf{W}_{0}^{l}\cdot\mathbf{h}_{d_{i}}^{l}+\sum_{r\in\mathcal{R}}\mathbf{W}_{r}^{l}\cdot(\frac{1}{c_{{d_{i}},r}}\sum_{j\in\mathcal{N}_{r}({d_{i}})}\mathbf{h}_{j}^{l})),$ (7)
$\mathbf{W}_{r}^{l}=I+\Theta_{r}^{l},$ (8)
@@ 4.3.2 Refinement of Point-to-Point Relationship
4.3.3 Aggregation of Entity Information
@@
$\footnotesize\begin{gathered}\mathbf{h}_{\mathcal{D}_{t}}=\text{Agg}(\mathbf{h}^{r}_{d},\mathbf{h}^{e}_{d}),\mathbf{h}_{\mathcal{P}_{t}}=\text{Agg}(\mathbf{h}^{r}_{p},\mathbf{h}^{e}_{p}),\\ \mathbf{h}_{\mathcal{M}_{t-1}}=\text{Agg}(\mathbf{h}^{r}_{m},\mathbf{h}^{e}_{m}),\end{gathered}$ (11)
$\mathbf{h}_{v_{t}}=[\mathbf{h}_{\mathcal{D}_{t}}||\mathbf{h}_{\mathcal{P}_{t}}||\mathbf{h}_{\mathcal{M}_{t-1}}].$ (12)
4.4 Medication Recommendation
@@
This module merges clinical visit information for predictions. Using multiple Gated Recurrent Units (GRUs) to capture temporal dependencies among visits, we generate the output of GRUs $\mathbf{o}_{v_{t}}$ and patient embedding $\mathbf{h}_{\mathcal{H}}$. Then, a Multi-Layer Perceptron (MLP) layer with activation function $\sigma$ converts this representation into medication scores. Medication $\hat{m}_{i}$ with a score exceeding $\delta$ is chosen as a result.
$\mathbf{h}_{\mathcal{H}}=\mathbf{o}_{v_{t}}=\text{GRU}(\mathbf{o}_{v_{t-1}},\mathbf{h}_{v_{t}}),$ (13)
$score=\sigma(\text{MLP}(\mathbf{h}_{\mathcal{H}})),$ (14)
$\hat{m}_{i}=\begin{cases}1,&score_{m_{i}}\geq\delta\\ 0,&score_{m_{i}}<\delta\end{cases}.$ (15)
4.5 Model Training and Inference
$\mathcal{L}_{bce}=-\sum_{i=1}^{|\mathcal{M}|}{m_{i}}\log({\hat{m}_{i}})+(1-{m_{i}})\log(1-{\hat{m}_{i}}),$ (16)
$\mathcal{L}_{multi}=\sum_{i,j:{m_{i}}=1,{m_{j}}=0}\frac{\max(0,1-({\hat{m}_{i}}-{\hat{m}_{j}}))}{|\mathcal{M}|},$ (17)
$\mathcal{L}_{ddi}=\sum_{i=1}^{|\mathcal{M}|}\sum_{j=1}^{|\mathcal{M}|}\mathbf{M}^{ddi}_{ij}\cdot{\hat{m}_{i}}\cdot{\hat{m}_{j}},$ (18)
We adopt a loss aggregation approach that aligns with the methodology used in prior work Yang et al. (2023).
$\footnotesize\mathcal{L}=\alpha(\beta\mathcal{L}_{bce}+(1-\beta)\mathcal{L}_{multi})+(1-\alpha)\mathcal{L}_{ddi},$ (19)
$\alpha=\begin{cases}1,&\text{DDI rate}\leq\gamma\\ \max\{0,1-\frac{\text{DDI rate}-\gamma}{kp}\},&\text{DDI rate}>\gamma\end{cases},$ (20)
4.6 Discussion on Causality
Compared to co-occurrence-based approaches, causality-based methods can offer more nuanced and clear relationships between medical entities, thereby enhancing personalized patient representations, as detailed in subsections 5.5 and 5.6.
5 Experiments
To validate the effectiveness of DisMed, we conduct experiments on real-world medical datasets.
5.1 Experiments Setup
Item MIMIC-III MIMIC-IV
# patients 6,350 60,125
# clinical events 15,032 156,810
# diseases 1,958 2,000
# procedures 1,430 1,500
# medications 131 131
avg. # of visits 2.37 2.61
avg. # of medications 11.44 6.66
Table 1: Statistics of the datasets.
5.1.1 Datasets and Metrics
We verify the model performance on the MIMIC-III Johnson et al. (2016) and MIMIC-IV Johnson et al. (2023). Our methodology for data processing and evaluation mirrored that of Yang et al. (2021b), encompassing: 1) Data preprocessing, with outcomes detailed in Table 1; 2) Division of datasets into training, validation, and test sets in a 4/6-1/6-1/6 ratio; 3) Utilization of Jaccard, F1-score, and PRAUC as accuracy metrics, and DDI-rate as a safety metric; 4) A sampling approach involving the extraction of test data from the test set via ten rounds of bootstrap sampling.
5.1.2 Baselines
LR is an instance-based classifier, that incorporates $L2$ regularization. ECC Read et al. (2011), an RS designed for multi-label classification. RETAIN Choi et al. (2016b) uses reverse time attention. LEAP Zhang et al. (2017) applies reinforcement learning for label dependencies and DDI. GAMENet Shang et al. (2019) utilizes a graph-augmented memory network for safe medication combinations. SafeDrug Yang et al. (2021b) focuses on DDI and leverages molecular structures. MICRON Yang et al. (2021a) predicts medication changes using residual networks. COGNet Wu et al. (2022) generates medication sets through a copy-and-predict mechanism. MoleRec Yang et al. (2023) focuses on the representation of molecular substructures.
5.1.3 Implementation Details
In our study, we configure embedding dimensions at $dim=64$ with a dropout rate of 0.5. The RGCN architecture consists of a 2-layer graph neural network segmented into 5 groups. For GRU and MLP, a single-layer network is employed, featuring a 64-dimensional hidden layer and ReLU activation. The recommended threshold $\delta=0.5$. The loss function parameters are set with $\beta=0.95$, $kp=0.05$, and an acceptance rate $\gamma=0.06$. Parameter training employs the Adam Optimizer with a learning rate of $lr=0.0005$, integrating $L2$ regularization during the training process with a regularization coefficient of 0.005. All experiments are conducted on an Ubuntu machine with 30GB memory, 12 CPUs, and a 24GB NVIDIA RTX3090 GPU.
5.2 Performance Comparison
@@
DisMed 0.5389 0.0709 0.6916 0.7826 20.5419 0.4899 0.0677 0.6412 0.7338 14.4295
@@
Table 2: The performance of each model on the test set regarding accuracy and safety. The best and the runner-up results are highlighted in bold and underlined respectively under t-tests, at the level of 95% confidence level.
@@
To demonstrate the effectiveness of DisMed, we conducted comparative experiments with several state-of-the-art baselines, and the results are shown in Table 2. In the realm of traditional machine learning-based methods, both LR and ECC demonstrate sub-optimal performance in terms of accuracy and safety. Despite employing deep learning strategies, LEAP fails to consider patients’ longitudinal historical visit information, resulting in no significant improvement in outcomes. While RETAIN and GAMENet effectively enhance accuracy, they overlook the consideration of DDI issues. In recent methodologies, SafeDrug introduces the significance of molecules, alleviating the mutual constraints between accuracy and safety. MICRON and COGNet approach the task as a translation, utilizing residual networks and transformers to further enhance effectiveness. However, they still exhibit certain deficiencies in terms of safety. MoleRec improves the accuracy of recommendations based on the learning representation of molecular substructures.
@@
DisMed introduces a condition-centric framework, recommending personalized medication combinations based on causality. Experimental results demonstrate a significant enhancement of our proposed method compared to other baselines, validating the effectiveness of DisMed.
5.3 Ablation Study
Figure 3: The performance of multiple variants of DisMed.
To verify the effectiveness of our proposed innovations, we evaluate three variant models on MIMIC-III. DisMed $w/o$ T omits the causal effects and the refinement of the point-to-point relationships between diseases/procedures and medications. DisMed $w/o$ P excludes pathological relationships, thus not identifying disease differences. DisMed $w/o$ T+P represents the model concurrent removal of both modules.
As illustrated in Figure 3, the result indicates that compared to DisMed $w/o$ T+P, both DisMed $w/o$ T and DisMed $w/o$ P show significant improvements in accuracy. This suggests that each of these two modules independently enhances the personalization of patient representations and boosts the accuracy of the outcomes. Notably, DisMed $w/o$ P exhibits a significantly higher DDI rate. This is attributed to our modeling of inter-medication relationships during the condition analysis stage, where we learn about the causal interactions between medications, thereby enhancing the safety of the outcomes. The higher performance of DisMed indicates that the two modules have a synergistic effect when used together, leading to a more pronounced enhancement. Through the construction of these variant models, we can verify the contributions of the different innovation points to our model.
5.4 Parameters Analysis
Edge type 4 5 6 7
@@ 5.4 Parameters Analysis
Table 3: The impact of three hyperparameters on the model.
In this subsection, we investigate the impact of several hyperparameters on MIMIC-III. These parameters include the number of types of edges in the bipartite graph, the number of layers in the RGCN, and the embedding dimensions for each medical entity. We select four neighboring configurations and utilize the Jaccard and DDI rate to represent the accuracy and safety of the results, as illustrated in Figure 3.

### v2

@@ 4.3.1. Homomorphic Relationship Learning
First, we extract relationships from the causal graph $G_{v_{t}}$ that exclusively involves entities from the current visit and construct three isomorphic causal graphs: $G^{d}_{v_{t}}$, $G^{p}_{v_{t}}$, and $G^{m}_{v_{t-1}}$. And then we categorize the patient’s disease/procedure, as shown in Figure 4. taking the example of diseases, for the learning of differences between diseases in a clinical visit, we categorize the diseases $\mathcal{D}_{t}$ into four types $\mathcal{D}^{j}_{t}$ based on $G^{d}_{v_{t}}$: 1) Causal diseases $\mathcal{D}^{1}_{t}$, which initiate other diseases and represent the primary ailment; 2) Effect diseases $\mathcal{D}^{2}_{t}$, influenced by other diseases and indicative of secondary symptoms; 3) Middle diseases $\mathcal{D}^{3}_{t}$, both influencing and influenced by other diseases; 4) Independent diseases $\mathcal{D}^{4}_{t}$, existing without direct causal links to other diseases.
Since each patient has a unique disease history and record, the same disease may occupy different positions in the causal graphs $G^{d}_{v_{t}}$ of different patients, leading to classification into different disease groups. The DSA mechanism learns the dynamic changes of disease groups in different health states, effectively capturing the personalized differences of diseases when facing different patients. It updates the representations of each entity based on these groupings to enhance the expressiveness of key entities. The specific formula is as follows:
(4) $\mathcal{D}^{j}_{t}=\text{Classify}(d_{i},G^{d}_{v_{t}}),$
(5) $\mathbf{h}^{r}_{d_{i}}=\mathbf{h}_{d_{i}}\cdot w^{j}_{t},$
(6) $w^{j}_{t}=\frac{\exp(\mathbf{W}\cdot\mathbf{h}^{D^{j}}_{t}+b)}{\sum_{k=1}^{4}\exp(\mathbf{W}\cdot\mathbf{h}^{D^{k}}_{t}+b)},$
@@
Figure 4. Using the classifier within DSA, we analyze the causal position of each node in the causal graph $G^{d}_{v_{t}}$ to gauge its influence on the patient’s current state. For instance, if $d_{3}$ is identified as a causal disease, it is classified as ${D}^{1}_{t}$.
[…]
4.3.2. Heterogeneous Relationship Learning
As shown in the right half of the Health State Learning part of Figure 2. At this stage, we update the representations of medical entities based on the correlations between diseases/procedures and medications obtained through causal estimation in relationship mining. This process aims to enhance the accuracy of entity representations.
We utilize the causal effect matrices $\mathbf{M}^{dm}$ and $\mathbf{M}^{pm}$ to construct disease-medication and procedure-medication complete bipartite graphs $G^{dm}_{v_{t}}$ and $G^{pm}_{v_{t}}$ for single visit.
Taking $G^{dm}_{v_{t}}$ as an example, the nodes encompass both $\mathcal{D}_{t}$ and $\mathcal{M}_{t-1}$. By utilizing Relational Graph Convolutional Networks (RGCN) (Schlichtkrull et al., 2018), we categorize edges corresponding to relationships with similar causal effects into the same type. Subsequently, we establish point-to-point relationships between diseases and medications. The representation of $d_{i}$ at layer $l$ is then updated on the point-to-point relationship using the following formula:
(7) $\mathbf{h}_{d_{i}}^{l+1}=\sigma(\mathbf{W}_{0}^{l}\cdot\mathbf{h}_{d_{i}}^{l}+\sum_{r\in\mathcal{R}}\mathbf{W}_{r}^{l}\cdot(\frac{1}{c_{{d_{i}},r}}\sum_{j\in\mathcal{N}_{r}({d_{i}})}\mathbf{h}_{j}^{l})),$
(8) $\mathbf{W}_{r}^{l}=I+\Theta_{r}^{l},$
@@ 4.3.2. Heterogeneous Relationship Learning
4.3.3. Aggregation of Entity Information
@@
(11) $\begin{gathered}\mathbf{h}_{\mathcal{D}_{t}}=\text{Agg}(\mathbf{h}^{r}_{d},\mathbf{h}^{e}_{d}),\mathbf{h}_{\mathcal{P}_{t}}=\text{Agg}(\mathbf{h}^{r}_{p},\mathbf{h}^{e}_{p}),\\ \mathbf{h}_{\mathcal{M}_{t-1}}=\text{Agg}(\mathbf{h}^{r}_{m},\mathbf{h}^{e}_{m}),\end{gathered}$
(12) $\mathbf{h}_{v_{t}}=[\mathbf{h}_{\mathcal{D}_{t}}||\mathbf{h}_{\mathcal{P}_{t}}||\mathbf{h}_{\mathcal{M}_{t-1}}].$
4.4. Medication Recommendation
@@
To deep mine personalized patient representations, we consider the patient’s medical history. We represent each medical record and integrate this information to recommend medication combinations. Using multiple GRUs to capture temporal dependencies among visits, we generate the GRU output $\mathbf{o}_{v_{t}}$ and patient embedding $\mathbf{h}_{\mathcal{H}}$. An MLP layer with activation function $\sigma$ converts this representation into medication scores. Medication $\hat{m}_{i}$ with a score exceeding $\delta$ will be chosen.
(13) $\mathbf{h}_{\mathcal{H}}=\mathbf{o}_{v_{t}}=\text{GRU}(\mathbf{o}_{v_{t-1}},\mathbf{h}_{v_{t}}),$
(14) $score=\sigma(\text{MLP}(\mathbf{h}_{\mathcal{H}})),$
(15) $\hat{m}_{i}=\begin{cases}1,&score_{m_{i}}\geq\delta\\ 0,&score_{m_{i}}<\delta\end{cases}.$
Table 1. The performance of each model on the test set regarding accuracy and safety. The best and the runner-up results are highlighted in bold and underlined respectively under t-tests, at the level of 95% confidence level.
@@
CausalMed 0.5389 0.0709 0.6916 0.7826 20.5419 0.4899 0.0677 0.6412 0.7338 14.4295
4.5. Model Training and Inference
(16) $\mathcal{L}_{bce}=-\sum_{i=1}^{|\mathcal{M}|}{m_{i}}\log({\hat{m}_{i}})+(1-{m_{i}})\log(1-{\hat{m}_{i}}),$
(17) $\mathcal{L}_{multi}=\sum_{i,j:{m_{i}}=1,{m_{j}}=0}\frac{\max(0,1-({\hat{m}_{i}}-{\hat{m}_{j}}))}{|\mathcal{M}|},$
(18) $\mathcal{L}_{ddi}=\sum_{i=1}^{|\mathcal{M}|}\sum_{j=1}^{|\mathcal{M}|}\mathbf{M}^{ddi}_{ij}\cdot{\hat{m}_{i}}\cdot{\hat{m}_{j}}.$
We adopt a loss aggregation approach that aligns with the methodology used in prior work (Yang et al., 2023).
(19) $\mathcal{L}=\alpha(\beta\mathcal{L}_{bce}+(1-\beta)\mathcal{L}_{multi})+(1-\alpha)\mathcal{L}_{ddi},$
(20) $\alpha=\begin{cases}1,&\text{DDI rate}\leq\gamma\\ \max\{0,1-\frac{\text{DDI rate}-\gamma}{kp}\},&\text{DDI rate}>\gamma\end{cases},$
5. Experiments
In this section, we conduct extensive experiments aimed at answering the following key research questions (RQ):
  • RQ1: Does CausalMed provide more accurate and safe recommendations than state-of-the-art medication recommendation systems?
  • RQ2: Can the two core modules proposed in this paper enhance the effectiveness of recommendations?
  • RQ3: Are the causal relationships used by CausalMed more appropriate for medication recommendations than the co-occurrence relationships used in previous medication recommendations?
  • RQ4: Why CausalMed can generate more personalized patient representations based on causality?
  • RQ5: Can the representation process of CausalMed be made transparent, and are its results interpretable?
5.1. Experiments Setup
Table 2. Statistics of the datasets.
Item MIMIC-III MIMIC-IV
# patients 6,350 60,125
# clinical events 15,032 156,810
# diseases 1,958 2,000
# procedures 1,430 1,500
# medications 131 131
avg. # of visits 2.37 2.61
avg. # of medications 11.44 6.66
5.1.1. Datasets
We verify the model performance on the MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023). Our methodology for data processing and evaluation mirrored that of (Yang et al., 2021b), encompassing: 1) Data preprocessing, with outcomes detailed in Table 2; 2) Division of datasets into training, validation, and test sets in a 4/6-1/6-1/6 ratio; 3) A sampling approach involving the extraction of test data from the test set via ten rounds of bootstrap sampling.
5.1.2. Evaluation Metrics
1)Jaccard: measures the overlap between predicted and actual sets, indicating similarity; 2) DDI-rate: Assesses safety by indicating the potential for harmful medication interactions. A lower value signifies a safer medication combination; 3)F1-Score: Balances precision and recall, a generic evaluation metric for recommendation tasks; 4)Precision-Recall AUC (PRAUC): Represents the model’s ability to distinguish between classes across various thresholds. 5)Avg.#Med: Reflects the average number of medications recommended per clinical visit and is used for reference only, not indicative of the recommendation system’s overall quality.
@@ 5.1.2. Evaluation Metrics
Table 3. The performance of multiple variants of CausalMed on evaluation metrics. The best and the runner-up results are highlighted in bold and underlined respectively under t-tests, at the level of 95% confidence level.
Model MIMIC-III MIMIC-IV
[…]
CausalMed w/o T 0.5369 0.0734 0.6900 0.7824 0.4878 0.0702 0.6405 0.7330
CausalMed w/o P 0.5324 0.0731 0.6860 0.7745 0.4838 0.0695 0.6355 0.7258
CausalMed w/o T+P 0.5339 0.0740 0.6873 0.7788 0.4847 0.0711 0.6373 0.7309
CausalMed 0.5389 0.0709 0.6916 0.7826 0.4899 0.0677 0.6412 0.7338
5.1.3. Baselines
LR (Indra et al., 2016) is a classifier for medication recommendation but doesn’t consider historical impacts on predictions. ECC (Read et al., 2011) improves multi-label classification by splitting it into multiple binary problems. RETAIN (Choi et al., 2016b) uses a reverse time attention mechanism to highlight influential historical events in sequential medical data. LEAP (Zhang et al., 2017) uses reinforcement learning to optimize medication strategies by predicting label dependencies and DDI. GAMENet (Shang et al., 2019) employs a graph-enhanced memory network to enhance the safety and efficacy of medication combinations. SafeDrug (Yang et al., 2021b) predicts DDI by analyzing structural similarities between medication molecules. MICRON (Yang et al., 2021a) uses a residual network to personalize medication recommendations by updating the latest combinations. COGNet (Wu et al., 2022) generates medication sets using cloning and prediction mechanisms, reintroducing significant past medications. MoleRec (Yang et al., 2023) analyzes molecular substructures to better predict medication properties and interactions.
5.1.4. Implementation Details
In our study, as shown in subsection 5.4, we configure embedding dimensions at $\text{dim}=64$. The RGCN architecture consists of a 2-layer graph neural network segmented into 5 groups. For GRU and MLP, a single-layer network with a 64-dimensional hidden layer and ReLU activation is employed. The parameters, consistent with previous work (Yang et al., 2023), include a dropout rate of 0.5, a threshold $\delta=0.5$, loss function parameters $\beta=0.95$, $kp=0.05$, and an acceptance rate $\gamma=0.06$. Training uses the Adam Optimizer with a learning rate of $lr=0.0005$ and $L2$ regularization with a coefficient of 0.005. Experiments are conducted on an Ubuntu machine with 30GB memory, 12 CPUs, and a 24GB NVIDIA RTX3090 GPU.
5.2. Performance Comparison (RQ1)
@@
To demonstrate the effectiveness of CausalMed, we conducted comparative experiments with several state-of-the-art baselines, and the results are shown in Table 1. In the realm of traditional machine learning-based methods, both LR and ECC demonstrate sub-optimal performance in terms of accuracy and safety. Despite employing deep learning strategies, LEAP fails to consider patients’ longitudinal historical visit information, resulting in no significant improvement in outcomes. Although RETAIN and GAMENet effectively improve accuracy, they ignore the consideration of DDI issues. SafeDrug introduces the importance of molecules and uses molecular characterization to identify potential DDI problems, greatly improving safety. However, it fails to balance accuracy and safety, performing poorly in terms of accuracy. MICRON proposes using the residual network to learn the difference in patient status between two consecutive visits. It does not directly prescribe medication but updates the last prescription. The results are greatly improved. COGNet treats this task as a translation task and uses Transformer to further improve efficiency. However, due to insufficient consideration of safety, there are still certain flaws. MoleRec conducts further research on molecules, using the substructure of molecules to learn medication representations and recommendations, improving in multiple dimensions.
@@
CausalMed introduces a health state-centric framework, recommending personalized medication combinations based on causality. Experimental results demonstrate a significant enhancement of our proposed method compared to other baselines, validating the effectiveness of CausalMed.
Table 4. The impact of three hyperparameters on the model. The best and the runner-up results are highlighted in bold and underlined respectively under t-tests, at the level of 95% confidence level. Parameters used in the final model are in bold.
parameter MIMIC-III MIMIC-IV
Edge type 4 5 6 7 4 5 6 7
@@
5.3. Ablation Study (RQ2&RQ3&RQ4)
To verify the effectiveness of our proposed innovations, we thoroughly evaluate three variant models. CausalMed $w/o$ T omits the causal effects and the refinement of the point-to-point relationships between diseases/procedures and medications. CausalMed $w/o$ P excludes pathological relationships, thereby not identifying disease differences. CausalMed $w/o$ T+P represents the model with the concurrent removal of both modules. Since the ablation versions originate from the same framework, we remove the Avg.#Med which only served as a reference.
As illustrated in Table 3, the result indicates that compared to CausalMed $w/o$ T+P, both CausalMed $w/o$ T and CausalMed $w/o$ P show significant improvements in accuracy. This indicates that CausalMed $w/o$ T can identify causal point-to-point associations between diseases/surgeries and medications, providing a more accurate relational basis through the graph network process. At the same time, CausalMed $w/o$ P can learn personalized representations based on the health state of the patient. Notably, CausalMed $w/o$ P exhibits a significantly higher DDI rate. This is attributed to our modeling of inter-medication relationships during the health state learning stage, where we learn about the causal interactions between medications, thereby enhancing the safety of the outcomes. The higher performance of CausalMed indicates that the two modules have a synergistic effect when used together, leading to a more pronounced enhancement. Through the construction of these variant models, we can effectively verify the contributions of the different innovation points to our overall model.
5.4. Parameters Analysis
In this subsection, we investigate the impact of several hyperparameters. These parameters include the number of types of edges in the bipartite graph, the number of layers in the RGCN, and the embedding dimensions for each medical entity. We select four neighboring configurations and utilize the Jaccard and DDI rate to represent the accuracy and safety of the results in Table 4.